Oak Ridge Waste Management has a subsidiary that disposes of hazardous waste and a subsidiary that collects and disposes of residential garbage. Information related to the two subsidiaries follows:

Hazardous Waste Residential Waste
Total Assets $15,232,900 $86,750,000
Noninterest-bearing current liabilities 3,272,200 13,028,200
Net income 1,889,900 6,642,300
Interest expense 1,450,100 7,990,100
Required Rate of return 10% 13%
Tax rate 40% 40%

Calculate ROI for both subsidiaries. (Round answers to 2 decimal places, e.g. 15.32%.)

Calculate EVA for both subsidiaries. Note that since no adjustments for accounting distortions are being made, EVA is equivalent to residual income. (Enter negative answers preceding either - sign, e.g. -45 or in parentheses, e.g. (45).)

Which subsidiary has added the most to shareholder value in the last year?

Based on the limited information, which subsidiary is the best candidate for expansion?
Answer #1
Return on income is calculated using the following expression : Economic value added is similar to theresidual income and is calculated as follows :
Return on income = Net operating profit after tax / Invested capital Economic value added = net operating income after tax - [Invested capital * cost of capital]
1) Net operating profit after tax = net income + Interest after tax EVA for hazardous waste = 2759960-(11960700*10%) $ 1,563,890
Hazardous waste = 18,89,900 + 14,50,100 (1-0.40) $          2,759,960 EVA for residential waste =   11436360-(73721800*13%) $ 1,852,526
Residential Waste = 6642300+7990100*(1-0.40) $        11,436,360
Now invested capital is calculated by deducting the non interest bearing liabilities from the total assets.
Invested capital = Total assets - Non interest bearing liabilities
Hazardous waste = 15232900-3272200 $        11,960,700
Residential Waste = 86750000 - 13028200 $        73,721,800
Return on Investment is Computed below :
Hazardous waste = 2759960/11960700 23.08%
Residential waste = 11436360/73721800 15.51%
c) Residential welfare subsidy has added the most shareholder value in the last year.
d) Hazardous waste subsidy is the best candidate for expansion because its return on investment is high and its WACC is also least